1. In 1975 tuition at Wattsomata University was $2,500 and the consumer price index was 80. In 2011 tuition was $12,000 and the price index was 320. Which of the following is correct?

            a.         Nominal and real tuition were both higher in 1975.

            b.         Nominal and real tuition were both higher in 2011.

            c.         Nominal tuition was higher in 1975, real tuition was higher in 2011.

            d.         Nominal tuition was higher in 2011, real tuition was higher in 1975.

Please show your calculations
Answer #1

Nominal Tuition fee in 1975 = $2,500

CPI = 80

As CPI is 20% less, we will multiply it by 1.2 to calculate real fee (our main motive is to take CPI = 100)

Real Tuition fee in 1975 = 2,500 * 1.2 = 3,000

Nominal Tuition fee in 2011 = $12,000

CPI = 320

Real Tuition fee in 2011 = (12,000 / 320) * 100 = 3,750

Thus, we can say that nominal and real tuition fee is higher in 2011 than 1975.

Option B is correct.